200.192.65.0 / 25
How many subnets does the chosen mask produce?
How many valid hosts per subnet are available?
What are the valid subnets?
What’s the broadcast address for each subnet?
What are the valid hosts in each subnet?
1
Expert's answer
2011-01-14T02:44:16-0500
It seems that this is the IP but not the subnet mask, you can find all the answers according to these rules:
255.255.255.XXX
Turn XXX into binary code, then
1. 2x – 2 = number of subnets. x is the number of the 1s. For example, in 11000000, the number of ones gives us 22 - 2 subnets. In this example, there are 2 subnets.
2. 2x – 2 = number of hosts per subnet. x is the number of the 0s. For example, in 11000000, the number of zeros gives us 26 - 2 hosts. In this example, there are 62 hosts per subnet.
3. 256 – subnet mask = block size, or base number. For example, (XXX=192) 256 – 192 = 64. 64 is the first subnet. The next subnet would be the base number itself, or 64 + 64 = 128, (the second subnet). You keep adding the base number to itself until you reach the value of the subnet mask, which is not a valid subnet because all subnet bits would be turned on (1s).
4. The broadcast address is all host bits turned on, which is the number immediately preceding the next subnet.
5. Valid hosts are the numbers between the subnets, minus all 0s and all 1s.